Order: Tylenol Elixir 240 mg p.o. for a child who weighs 21 kg
The recommended dosage is 1 tsp for weight of 24 to 35 lb or 1 1/2 tsp for weight of 36 to 47 lb.
Available:NDC 11673-130-26 childrens acetaminophen oral suspension 80 mg per 2 teaspoon (160 mg per 5 mL) fever reducer/pain reliever

How many Milliliters should the client receive?
Answer #1

ANSWER.

* Calculation of volume(ml) of drug the client receive.

       Formula

          Volume(ml) = Desired dose /Dose in hand * Quantity.

    Here

         Ordered dose = 240 mg for a 21 kg child.

       Convert kg to lb for calculation

                1 kg = 2.205 lb

         So

           weight of a 21 kg child in lb = 21 * 2.205

                                                       = 46.3 lb.

        Recommended dose for this child is 1 1/2 teaspoon.

        Desired dose (D) = 240 mg

      Dose in hand (H) = 240 mg ( 160 + 80 = 1 1/2 teaspoon)

              Quantity (Q) = 7.5 ml ( 1 1/2 teaspoon)

    As per above formula

        Volume(ml) drug should the client receive = D/H * Q

                                                                             = 240/240 * 7.5

                                                                             = 7.5 ml.